Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation:
s= c-1/3
If you wear a size 7 cap, what is your head circumference

Answer by stanbon(75887)

Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation:
s= (c-1)/3
If you wear a size 7 cap, what is your head circumference
------------------
7 = (c-1)/3
c-1 = 21
circumference = 22 inches
-----------------------------
cheers,
Stan H.